How much work must be done in blowing a spherical soap bubble 20 cm in diameter? The area of a sphere is 4πr² and the surface tension of the soap solution is 0.032 N/m.
1
Expert's answer
2020-11-27T10:09:16-0500

The work is


"W=\sigma\Delta A=\sigma(A_2-A_1),\\
W=4\pi\sigma(R^2-r^2)=\\
=4\cdot3.14\cdot0.032(0.1^2-0)=4.02\cdot10^{-3}\text{ J}."